OpenStudy (anonymous):

write the partial fraction decomposition for the rational expression. Check your result algebraically by combining the fractions. 1/(x^2 -1)

OpenStudy (anonymous):

\[ \large \frac{1}{x^2-1} \]First you factor it with difference of squares\[ \frac{1}{(x+1)(x-1)} \]Now we want this form:\[ \frac{1}{(x+1)(x-1)} = \frac{A}{x+1}+\frac{B}{x-1} \]We know that\[ A(x-1)+B(x+1)=1 \]Since it was used in order to get the numerator of 1.

OpenStudy (anonymous):

We get\[ Ax-A+Bx+B = 1 \]\[ (A+B)x+(B-A) = 1 \]

OpenStudy (anonymous):

So \[ (A+B)x+(B-A) = 0x+1 \]Gives us two equations:\[ A+B = 0\quad\quad B-A = 1 \]